# In your own words: computationally identifying interpretable themes from free-text survey data

### 27 Mar 2026 v1

@@ Appendix C Additional details on the computational framework
The prompt shown in Figure S5 was adapted to support annotation of identity-related free-text responses. To select an annotation model configuration, we evaluated several OpenAI model variants using both this identity-specific prompt and a baseline prompt drawn from prior work on SAE interpretation [hypothesaes]. Specifically, we compared GPT-4.1 and GPT-4.1-mini across model–prompt combinations and assessed performance relative to human annotations using three criteria: overall annotation agreement, consistency of agreement across race, gender, and sexual orientation themes, and how closely the model matches humans in how often it assigns a theme (positive annotation rate). Based on these criteria, we selected GPT-4.1-mini with the identity-specific prompt for use in our final analyses.
@@ D.1 Detailed presentation of free-text themes
| 15 | uses single-word or very short descriptions of race/ethnicity (1–3 words) | 0.69 |
@@
| 20 | mentions parentage explicitly (e.g., ’my mom is X and my dad is Y’) | 0.86 |
@@
| 32 | explicitly self-describes as non-binary or uses the term ’non-binary’ to describe their gender identity | 0.98 |
@@ D.2 Comparison of computationally identified free-text themes to standardized categories
For gender, we construct categories based on the cross-tabulation of current gender identity (e.g., “Man”, “Woman”, “Some other way”) [amaya2020adapting] and transgender status (“Yes”, “No”, “I prefer not to answer”). Respondents who select “I prefer not to answer” for either question are grouped into a single Prefer not to answer category. This yields seven mutually exclusive gender categories: Cisgender Man, Cisgender Woman, Cisgender Other, Transgender Man, Transgender Woman, Transgender Other, and Prefer not to answer.
For sexual orientation, we construct groupings by drawing on terminology and definitions from The Trevor Project and the Human Rights Campaign [TrevorProject2021Gay, TrevorProjectBisexuality, HRC2023Queer], while recognizing that collapsing distinct identities into shared categories (e.g., Bisexual and Pansexual) may not reflect how all individuals understand their own identities. Responses of “Gay” and “Lesbian” are grouped into the analytical category Gay or Lesbian [TrevorProject2021Gay], and “Bisexual” and “Pansexual” are grouped as Bisexual and/or Pansexual [TrevorProjectBisexuality]. Respondents who selected “Queer” alongside another minority identity are categorized under the more specific label, recognizing that “Queer” often serves as an umbrella term [HRC2023Queer]. Respondents who select two or more options after these categorizations are assigned to the Multiple Identities category. The final set of 11 categories includes Asexual or aromantic, Bisexual and/or Pansexual, Demisexual, Gay or Lesbian, I prefer not to answer, Multiple Identities, Other, Queer, Questioning, Sexually fluid, and Straight or heterosexual.
@@
| mentions parentage explicitly (e.g., ’my mom is X and my dad is Y’) | 0.08 |
@@
| uses the word ’cisgender’ | 0.01 |
@@
| uses the phrase ’I am a female’ | 0.09 |
@@
| explicitly states ’I am a woman’ | 0.12 |
@@
| uses the exact word ’male’ | 0.41 |
@@
| explicitly self-describes as non-binary or uses the term ’non-binary’ to describe their gender identity | 0.48 |
@@
| mentions ’opposite sex’ or ’opposite gender’ | 0.05 |
| explicitly states attraction to ’men’ without additional context or qualifiers | 0.05 |
@@
| mentions the word ’straight’ | 0.21 |
| explicitly self-describes as straight using the exact word ’straight’ | 0.24 |
| explicitly uses the term ’heterosexual’ | 0.25 |
@@
| uses the term ’queer’ to describe their sexual orientation | 0.46 |
@@ D.3 Additional details on ‘Free-text themes reveal within-group heterogeneity’
Here we provide additional details supporting the analyses which examine whether free-text themes can help explain additional variation in key health and social measures. We report full regression results for all outcome–identity pairs and visualize the associations between free-text themes and life outcomes after controlling for standardized identity categories. Table S10 presents the complete nested model comparison results, while Figures S10–S12 show theme-level associations for race, gender, and sexual orientation, respectively.
@@ F.1 Hyperparameters
BERTopic (grootendorst2022bertopic, grootendorst2022bertopic). BERTopic produces a variable number of themes—each represented by a set of terms weighted by how unique they are to the cluster—and assigns each document to a single theme. We run BERTopic on the same pretrained OpenAI text embeddings used in the In Your Own Words framework (M.3), with UMAP for dimensionality reduction, HDBSCAN for clustering, and c-TF–IDF for theme representation. This aligns with prior work (e.g., [hypothesaes, pham-etal-2024-topicgpt]). All components are left at their default settings except for the HDBSCAN minimum cluster size hyperparameter, which controls theme granularity.777https://github.com/MaartenGr/BERTopic We set the minimum cluster size to 7, which generates a comparable number of themes (23–29 per identity axis) to the number of free-text themes retained by our computational framework after filtering (26–28 per axis). After fitting BERTopic, we fine-tune the themes with an LLM label, using BERTopic’s built-in LLM representation function.
@@ F.2 Baseline Results
Without early stopping, the resulting theme structure is highly uneven across identity axes. For race, TopicGPT produces only two themes (“Cultural Identity” and “Agriculture”), with “Cultural Identity” assigned to 99% of responses, indicating that nearly the entire dataset is collapsed into a single, overly broad theme. For gender, TopicGPT generates eleven themes, but the dominant theme (“Gender Identity”) is assigned to 98% of responses, again resulting in minimal differentiation across responses. In contrast, for sexual orientation, TopicGPT produces over 50 themes, many of which occur infrequently, overlap in content, and redundantly encode similar identity labels (e.g., “Sexuality”, “Sexual Orientation”). Although TopicGPT’s subtheme generation can recover specific identity categories within the highest-frequency themes, this added hierarchical structure does not fix its core issue in our setting: the theme–subtheme hierarchy is not consistent across the race, gender, and sexual orientation responses. Instead, it introduces additional redundancy, with some concepts appearing as both themes and subthemes (e.g., “Pansexuality” appears as a subtheme under “Sexuality”, but also appears as a redundant standalone theme with 6.8% of responses). TopicGPT also produces themes that are weakly aligned with the free-text responses (e.g.,“Agriculture” for race; “Social Activities” and “Outdoor Activities” for gender). Together, these issues make TopicGPT difficult to use for systematic analysis or to adapt to specific analytic goals. In contrast, our computational framework produces a stable set of interpretable free-text themes, with explicit control over their granularity. This supports downstream analysis without the uneven coverage and sensitivity to response ordering we observe in TopicGPT.
@@ Appendix G Varying the number of computationally identified free-text themes
Acknowledgements The authors thank Serina Chang, Evan Dong, Frauke Kreuter, Hima Lakkaraju, Rajiv Movva, Katie Piner, Divya Shanmugam, David Shin, and participants at the Pew Research Center research methods seminar for helpful comments. This work was supported by an NSF Graduate Research Fellowship, a Google Research Scholar award, an AI2050 Early Career Fellowship, NSF CAREER #2142419, a CIFAR Azrieli Global scholarship, a gift to the LinkedIn-Cornell Bowers CIS Strategic Partnership, OpenAI API credits, the Survival and Flourishing Fund, Coefficient Giving, and the Zhang Family Endowed professorship.

### 03 Apr 2026 v2

@@ Appendix C Additional details on the computational framework
The prompt shown in Figure S5 was adapted to support annotation of identity-related free-text responses. To select an annotation model configuration, we evaluated several OpenAI model variants using both this identity-specific prompt and a baseline prompt drawn from prior work on SAE interpretation [hypothesaes]. Specifically, we compared GPT-4.1 and GPT-4.1-mini across model-prompt combinations and assessed performance relative to human annotations using three criteria: overall annotation agreement, consistency of agreement across race, gender, and sexual orientation themes, and how closely the model matches humans in how often it assigns a theme (positive annotation rate). Based on these criteria, we selected GPT-4.1-mini with the identity-specific prompt for use in our final analyses.
@@ D.1 Detailed presentation of free-text themes
| 15 | uses single-word or very short descriptions of race/ethnicity (1-3 words) | 0.69 |
@@
| 20 | mentions parentage explicitly (e.g., ‘my mom is X and my dad is Y’) | 0.86 |
@@
| 32 | explicitly self-describes as non-binary or uses the term ‘non-binary’ to describe their gender identity | 0.98 |
@@ D.2 Comparison of computationally identified free-text themes to standardized categories
For gender, we construct categories based on the cross-tabulation of current gender identity (e.g., “Man”, “Woman”, “Some other way”) [amaya2020adapting] and transgender status (“Yes”, “No”, “Prefer not to answer”). Respondents who select “Prefer not to answer” for transgender status and any gender identity are grouped into a single Prefer not to answer category. This yields seven mutually exclusive gender categories: Cisgender, Man; Cisgender, Woman; Cisgender, Some other way; Transgender, Man; Transgender, Woman; Transgender, Some other way; and Prefer not to answer.
For sexual orientation, we construct groupings by drawing on terminology and definitions from The Trevor Project and the Human Rights Campaign [TrevorProject2021Gay, TrevorProjectBisexuality, HRC2023Queer], while recognizing that collapsing distinct identities into shared categories (e.g., Bisexual and Pansexual) may not reflect how all individuals understand their own identities. Responses of “Gay” and “Lesbian” are grouped into the analytical category Gay or Lesbian [TrevorProject2021Gay], and “Bisexual” and “Pansexual” are grouped as Bisexual and/or Pansexual [TrevorProjectBisexuality]. Respondents who selected “Queer” alongside another minority identity are categorized under the more specific label, recognizing that “Queer” often serves as an umbrella term [HRC2023Queer]. Respondents who select two or more options after these categorizations are assigned to the Multiple Identities category. The final set of 11 categories includes Asexual or aromantic, Bisexual and/or Pansexual, Demisexual, Gay or Lesbian, Multiple Identities, Other sexual identity or orientation, Prefer not to answer, Queer, Questioning, Sexually fluid, and Straight or heterosexual.
@@
| mentions parentage explicitly (e.g., ‘my mom is X and my dad is Y’) | 0.08 |
@@
| uses the word ‘cisgender’ | 0.01 |
@@
| uses the phrase ‘I am a female’ | 0.09 |
@@
| explicitly states ‘I am a woman’ | 0.12 |
@@
| uses the exact word ‘male’ | 0.41 |
@@
| explicitly self-describes as non-binary or uses the term ‘non-binary’ to describe their gender identity | 0.48 |
@@
| mentions ‘opposite sex’ or ‘opposite gender’ | 0.05 |
| explicitly states attraction to ‘men’ without additional context or qualifiers | 0.05 |
@@
| mentions the word ‘straight’ | 0.21 |
| explicitly self-describes as straight using the exact word ‘straight’ | 0.24 |
| explicitly uses the term ‘heterosexual’ | 0.25 |
@@
| uses the term ‘queer’ to describe their sexual orientation | 0.46 |
@@ D.3 Additional details on ‘Free-text themes reveal within-group heterogeneity’
Here we provide additional details supporting the analyses which examine whether free-text themes can help explain additional variation in key health and social measures. We report full regression results for all outcome-identity pairs and visualize the associations between free-text themes and life outcomes after controlling for standardized identity categories. Table S10 presents the complete nested model comparison results, while Figures S10-S12 show theme-level associations for race, gender, and sexual orientation, respectively.
@@ F.1 Hyperparameters
BERTopic (grootendorst2022bertopic, grootendorst2022bertopic). BERTopic produces a variable number of themes—each represented by a set of terms weighted by how unique they are to the cluster—and assigns each document to a single theme. We run BERTopic on the same pretrained OpenAI text embeddings used in the In Your Own Words framework (M.3), with UMAP for dimensionality reduction, HDBSCAN for clustering, and c-TF-IDF for theme representation. This aligns with prior work (e.g., [hypothesaes, pham-etal-2024-topicgpt]). All components are left at their default settings except for the HDBSCAN minimum cluster size hyperparameter, which controls theme granularity.777https://github.com/MaartenGr/BERTopic We set the minimum cluster size to 7, which generates a comparable number of themes (23-29 per identity axis) to the number of free-text themes retained by our computational framework after filtering (26-28 per axis). After fitting BERTopic, we fine-tune the themes with an LLM label, using BERTopic’s built-in LLM representation function.
@@ F.2 Baseline Results
Without early stopping, the resulting theme structure is highly uneven across identity axes. For race, TopicGPT produces only two themes (“Cultural Identity” and “Agriculture”), with “Cultural Identity” assigned to 99% of responses, indicating that nearly the entire dataset is collapsed into a single, overly broad theme. For gender, TopicGPT generates eleven themes, but the dominant theme (“Gender Identity”) is assigned to 98% of responses, again resulting in minimal differentiation across responses. In contrast, for sexual orientation, TopicGPT produces over 50 themes, many of which occur infrequently, overlap in content, and redundantly encode similar identity labels (e.g., “Sexuality”, “Sexual Orientation”). Although TopicGPT’s subtheme generation can recover specific identity categories within the highest-frequency themes, this added hierarchical structure does not fix its core issue in our setting: the theme-subtheme hierarchy is not consistent across the race, gender, and sexual orientation responses. Instead, it introduces additional redundancy, with some concepts appearing as both themes and subthemes (e.g., “Pansexuality” appears as a subtheme under “Sexuality”, but also appears as a redundant standalone theme with 6.8% of responses). TopicGPT also produces themes that are weakly aligned with the free-text responses (e.g.,“Agriculture” for race; “Social Activities” and “Outdoor Activities” for gender). Together, these issues make TopicGPT difficult to use for systematic analysis or to adapt to specific analytic goals. In contrast, our computational framework produces a stable set of interpretable free-text themes, with explicit control over their granularity. This supports downstream analysis without the uneven coverage and sensitivity to response ordering we observe in TopicGPT.
@@ Appendix G Varying the number of computationally identified free-text themes
Acknowledgements The authors thank Serina Chang, Evan Dong, Frauke Kreuter, Hima Lakkaraju, Rajiv Movva, Eni Mustafaraj, Katie Piner, Divya Shanmugam, David Shin, and participants at the Pew Research Center research methods seminar for helpful comments. This work was supported by an NSF Graduate Research Fellowship, a Google Research Scholar award, an AI2050 Early Career Fellowship, NSF CAREER #2142419, a CIFAR Azrieli Global scholarship, a gift to the LinkedIn-Cornell Bowers CIS Strategic Partnership, OpenAI API credits, the Survival and Flourishing Fund, Coefficient Giving, and the Zhang Family Endowed professorship.